OpenStudy (anonymous):

What is the solution set of |x + 3| - 2 = -5?

OpenStudy (rock_mit182):

\[|x+3|=-3\] adding up both sides of the equation by 2

myininaya (myininaya):

@MilenaSaeger Please do not just give the answer. (Also this is not the right answer)

OpenStudy (anonymous):

wouls the answer be no solution?

OpenStudy (rock_mit182):

no you have to solve the absolute fvalue now

myininaya (myininaya):

@makaylarrrrrr and why did you reach that conclusion? not saying you are wrong

OpenStudy (anonymous):

because I plugged 6 and 0 back into the problem and they weren't equal to -5

myininaya (myininaya):

Ok @makaylarrrrrr you are right to say no solution because the absolute value of something would leave you to either a positive or 0 answer but plugging in any value of x here: |x+3| would leave you to a positive or 0 answer so there cannot possibly be a solution to |x+3|=-3

myininaya (myininaya):

gj @makaylarrrrrr

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

And @MilenaSaeger plugging in your answers as @makaylarrrrrr did you will see they do not work what did you say 6 and 0 the equation was |x+3|-2=-5 plug in 6 we get: |6+3|-2 |9|-2 9-2 7 but 7 is not -5 so x=6 will definitely not work you also said 0 would work but when we replace x with 0 we get |0+3|-2 |3|-2 3-2 1 but 1 is not -5 so x=0 will definitely not work and again |of some number| will either be positive or zero (and never negative)